# Forgetting Order of Continual Learning: Examples That are Learned First are Forgotten Last

### v1

@@ Appendix B Correlations between learning speed and catastrophic forgetting in different datasets
In the main paper, we omitted the standard errors for visualization porpuses from Figs. 3.2,5,6 and Table 3.5. In both cases, these were usually very small and did not affect the qualitative results. For completeness, we report the omitted values in this section. The standard error for Figs. 3.2,5,6 can be found in Figs. C,15,16 respectively. A replication of Table 3.5, with the standard errors, can be found in Table C.
(a) CIFAR-100-2, $4\%$ buffer
(b) CIFAR-100-2, $40\%$ buffer
(c) TinyImagenet-2, $20\%$ buffer
Figure 14: Standard errors for Fig. 3.2. The error is taken over 10 repetitions in each experiment.
(d) (A)–>(B)
(e) (A)–>(C)
@@
Table 2: Same as Table 3.5, but with the addition of standard errors. Each entry is the mean test accuracy across all tasks of 10 networks trained using experience replay on CIFAR-100-2. Bold denotes the method with the highest accuracy.
CIFAR-10-2 CIFAR-100-2
Buffer size $4\%$ $40\%$ $4\%$ $40\%$
Random 87.67 $\pm$ 0.06 95.03 $\pm$ 0.05 69.03 $\pm$ 8.89 79.24 $\pm$ 1.82
Max entropy 84.44 $\pm$ 0.23 94.63 $\pm$ 0.06 64.27 $\pm$ 2.11 77.91 $\pm$ 0.91
IPM 85.9 $\pm$ 0.16 94.76 $\pm$ 0.13 64.16 $\pm$ 4.13 75.73 $\pm$ 1.43
GSS 85.28 $\pm$ 0.13 95.55 $\pm$ 0.08 64.15 $\pm$ 5.01 80.11 $\pm$ 0.81
Herding 88.41 $\pm$ 0.09 94.53 $\pm$ 0.12 71.07 $\pm$ 9.49 78.97 $\pm$ 0.1
Goldilocks 88.99 $\pm$ 0.03 96.03 $\pm$ 0.1 71.53 $\pm$ 7.31 80.62 $\pm$ 1.11
(a) CIFAR-100-5
(b) CIFAR-10-5
(c) TinyImagenet-10
Figure 17: Goldilocks vs. uniform sampling with various continual learning methods. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error. (a) CIFAR-100-5 with a buffer size of $5\%$. (b) CIFAR-10-5 with a buffer of size $5\%$ (c) TinyImagenet-10 with a buffer of size $10\%$. Goldilocks significantly improves performance across all methods and datasets.
@@
Figure 18: Comparing replay buffer compositions when training the same original task, followed by different subsequent tasks. Each entry in the matrices denotes the mean final test accuracy of $10$ networks, trained with experience replay on 2 tasks, with a buffer size of $40\%$. This Figure replicated the same experimental setup as Fig. 5. In (a) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100. In (b) we train the models on the 50 first classes of CIFAR-100, followed by a rotation classification task on the same examples (see text). In (c) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100, but with random labels. In all cases, the same buffer compositions of the first task give the same qualitative results, suggesting that the optimal composition of the buffer of the first task is independent of the subsequent task.
Appendix D Independency of the optimal buffer composition when changing the subsequent tasks
@@
In Section 3.3, we demonstrate that when selecting tasks A, B, and C from the same dataset, the optimal buffer composition of task A remains unaffected by tasks B or C. We observe consistent quantitative behavior across different $quick$ and $slow$ parameters. We further explore task variations by setting A as the first 50 classes of CIFAR-100, B as the last 50 classes of CIFAR-100, and C as a rotation classification [13] task. In the rotation classification task, examples from A are randomly rotated by angles of $\{90\degree,180\degree,270\degree\}$, with labels adjusted accordingly. We evaluate the performance of 10 networks trained continuously with different buffer compositions first on task A and then on task B (Fig.18(a)), and on task A followed by task C (rotation) (Fig.18(b)). Consistent with previous findings, the performance of various buffer compositions remains consistent, independent of the choice of B or C. This property enables us to utilize the rotation task for hyper-parameter grid search without requiring additional label data.
We further modify task C to comprise examples from the last 50 classes of CIFAR-100 but with random labels [57]. Training on random labels forces the network to memorize the examples, as no generalization is possible, eliminating any potential overlap between the subsequent tasks. The results of training on task A followed by the random label task are depicted in Fig. 18(c). Analogous to the rotation classification scenario, we observe that the same buffer compositions of task A remain effective, further suggesting that the optimal buffer composition of A does not depend on the subsequent task.
Appendix E Goldilocks with different continual learning methods
Here, we compare Goldilocks to uniform sampling with different continual learning algorithms. We tested several different algorithms on CIFAR-100-5, CIFAR-10-5, and TinyImagenet-10. The results, which extend Fig. 7, are plotted in Fig. 17.

### v2

@@ Appendix E Standard Errors
In the main paper, we omitted the standard errors for visualization porpuses from Figs. 4,5 and Table 3.2. In all cases, these were usually very small and did not affect the qualitative results. For completeness, we report the omitted values in this section. The standard error for Figs. 4,5 can be found in Figs. 22,23 respectively. The standard errors of Table 3.2 can be found in Table E.
(a) CIFAR-100-2, $1k$ buffer
(b) CIFAR-100-2, $10k$ buffer
(c) TinyImageNet-2, $1k$ buffer
@@
Table 4: Standard error of Table 3.2.
CIFAR-10-2 CIFAR-100-2 CIFAR-100-20 CIFAR-10-5 TinyImageNet-2
Buffer size $1k$ $10k$ $1k$ $10k$ $1k$ $10k$ $1k$ $10k$ $1k$ $10k$
Random 0.06 0.05 0.22 0.12 0.18 0.04 0.1 0.11 0.15 0.07
Max entropy 0.23 0.06 0.23 0.15 0.12 0.02 0.1 0.03 0.2 0.13
IPM 0.16 0.13 0.29 0.16 0.17 0.06 0.15 0.06 0.15 0.13
GSS 0.13 0.08 0.29 0.16 0.15 0.01 0.11 0.04 0.13 0.07
Herding 0.09 0.12 0.24 0.14 0.11 0.1 0.14 0.06 0.16 0.11
LARS 0.15 0.12 0.27 0.18 0.16 0.04 0.11 0.05 0.13 0.08
SBS-fix 0.08 0.1 0.25 0.11 0.08 0.07 0.13 0.11 0.19 0.1
SBS 0.03 0.1 0.22 0.1 0.1 0.02 0.1 0.12 0.16 0.1
@@
Figure 24: Comparing replay buffer compositions when training the same original task, followed by different subsequent tasks. Each entry in the matrices denotes the difference in final accuracy between each buffer composition and a buffer sampled uniformly at random. Each composition is repeated $10$ times. Networks are trained using experience replay on 2 tasks, with a buffer size of $10k$. This Figure replicated the same experimental setup as Fig. 5. In (a) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100. In (b) we train the models on the 50 first classes of CIFAR-100, followed by a rotation classification task on the same examples (see text). In (c) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100, but with random labels. In all cases, the same buffer compositions of the first task give the same qualitative results, suggesting that the optimal composition of the buffer of the first task is independent of the subsequent task.
Appendix F Independency of the Optimal Buffer Composition When Changing the Subsequent Tasks
@@
In Section 2.6, we demonstrate that when selecting tasks A, B, and C from the same dataset, the optimal buffer composition of task A remains unaffected by tasks B or C. We observe consistent quantitative behavior across different $quick$ and $slow$ parameters. We further explore task variations by setting A as the first 50 classes of CIFAR-100, B as the last 50 classes of CIFAR-100, and C as a rotation classification (Gidaris et al., 2018) task. In the rotation classification task, examples from A are randomly rotated by angles of $\{90\degree,180\degree,270\degree\}$, with labels adjusted accordingly. We evaluate the performance of 10 networks trained continuously with different buffer compositions first on task A and then on task B (Fig. 24(a)), and on task A followed by task C (rotation) (Fig. 24(b)). Consistent with previous findings, the performance of various buffer compositions remains consistent, independent of the choice of B or C. This property enables us to utilize the rotation task for hyper-parameter grid search without requiring additional label data.
We further modify task C to comprise examples from the last 50 classes of CIFAR-100 but with random labels (Zhang et al., 2021). Training on random labels forces the network to memorize the examples, as no generalization is possible, eliminating any potential overlap between the subsequent tasks. The results of training on task A followed by the random label task are depicted in Fig. 24(c). Analogous to the rotation classification scenario, we observe that the same buffer compositions of task A remain effective, further suggesting that the optimal buffer composition of A does not depend on the subsequent task.
Appendix G Relationship Between Other Scoring Methods and the Learning Speed
(a) learning speed
(b) uncertainty score
(c) loss score
(d) c-score
Figure 25: Comparison of the extended epoch-wise classification matrix for CIFAR-100-2, sorted by different example-level scoring functions (see App. G). (a) Learning speed (used throughout the paper), (b) Uncertainty measure from Bang et al. (2021), (c) Final loss, (d) c-score from Jiang et al. (2020). While all scores show some correlation with forgetting, as more complex examples tend to be forgotten more, learning speed shows the strongest, suggesting it is most suitable for our analysis.
To analyze catastrophic forgetting, we focus on its relationship with learning speed (Eq.1). While other metrics exist—such as uncertainty score (Bang et al., 2021), which evaluates classification consistency under augmentations, and c-score (Jiang et al., 2020), which measures expected accuracy on held-out instances – they incur higher computational costs. Using per-example loss as a score is computationally simpler but shows a weak correlation with catastrophic forgetting. As shown in Fig.25, examples sorted by uncertainty and c-score show a moderate correlation with forgetting, while learning speed provides the strongest correlation. Therefore, we adopt learning speed throughout this work for its computational efficiency and strong alignment with catastrophic forgetting.